OpenStudy (loveneenah):

What is the answer to this math problem, 8x+9=5x+6(7-4) plz hlep!!!

OpenStudy (morgan97lw):

x=3

OpenStudy (morgan97lw):

Here's the explanation, First work what you have in the parentheses. so, 8x+9=5x+6(7-4) 8x+9=5x+6(3) Multiply 8x+9=5x+18 Isolate the variable to one side by subtracting 9 from both sides. 8x=5x+9 Subtract 5x from both sides 3x=9 Finally divide, and your final answer is x=3

OpenStudy (ladycielphantomhive):

This should be in the math section.